(1) 42,100 (2) 2,950 (7) 16,850l(3) 5,450 (3) 30,300 (6) 12,650 9) 4,200 5) 22,750 (2) 2,950(8) 1,700 (1) 42,100 (8) 1,700 1019 PM 1/22/2019 匍闸
Answer #1
Unadjusted Trial balance
Debit $ Credit $
Cash 29050
Accounts receivable 5900
Supplies 1250
Equipment 30300
Accounts payable 12200
Common Stock 42100
Dividends 4200
Service revenue 22750
Operating expense 6350
Total 77050 77050
Note:
Cash account-Debit balance = (42100+16850-2950-5450-4650-12650-4200)= 29050
Accounts receivable-Debit balance =22750-16850 = 5900
supplies- Debit balance =2950-700= 1250
Equipment-Debit =30300
Accounts payable- Credit balance =24850-12650 = 12200
Common Sttock-Credit - 42100
Dividend-Debit = 4200
Service revenue- Credit = 22750
Operating expense-Debit = 6350